Im so confused on how to set this question up.. Any help will be GREAT!
If you are 2/7 mile from home and you can walk 3 5/7 mph, how long will it take you to get home?

Answer by nerdybill(7384)

If you are 2/7 mile from home and you can walk 3 5/7 mph, how long will it take you to get home?
start by converting 3 5/7 into an improper fraction:
3 + 5/7 = 21/7 + 5/7 = 26/7 mph
.
then apply distance formula:
d = rt
2/7 = (26/7)t
multiplying both sides by 7
2 = 26t
2/26 = t
.077 hours = t
or
4.6 minutes = t